40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Aug 30 17:31:56 2019, anchal, DailyProgress, TempCtrl, Vacuum Can Temperature Control Setpoint Step Test 
    Reply  Wed Sep 4 11:04:08 2019, anchal, DailyProgress, TempCtrl, Vacuum Can Temperature Control Setpoint Step Test - Results VacCanStepResponse.pdfVacCanStepResponseFit.pdf
       Reply  Wed Sep 4 11:49:02 2019, rana, DailyProgress, TempCtrl, Vacuum Can Temperature Control Setpoint Step Test - Results 
          Reply  Thu Sep 5 11:15:06 2019, anchal, DailyProgress, TempCtrl, Vacuum Can Temperature Control Setpoint Step Test - Results VacCanStepResponseTempAxis.pdf
Message ID: 2415     Entry time: Thu Sep 5 11:15:06 2019     In reply to: 2411
Author: anchal 
Type: DailyProgress 
Category: TempCtrl 
Subject: Vacuum Can Temperature Control Setpoint Step Test - Results 

I used the calibration of North and South NPRO control voltage to frequency change from CTN:1948.

Then using the following formula:

\frac{\Delta T}{\Delta V} = \frac{S[\frac{Hz}{V}]}{-\alpha [\frac{1}{K}]\frac{c[\frac{m}{s}]}{\lambda [m]}}

where S is the calibration slope for NPRO from volts to the frequency

\alpha is the coefficient of thermal expansion of fused silica, 5.5\times10^{-7} K^{-1}

c is the speed of light 299792458 m/s

The offset was adjusted so that at t=0, the cavity temperatures is the same as the vacuum can in-loop temperature sensor reading at t=0.

This gave for the South Cavity -22.456 K/V and for the North Cavity -23.489 K/V

The South Cavity temperature changed 0.511 K, 0.539 K, and 0.565 K in the first step on the three days. Mean response to the first step is 0.538 K.

The North Cavity temperature changed 0.520 K, 0.574 K, and 0.639 K in the first step on the three days. Mean response to the first step is 0.578 K.

The South Cavity temperature changed -0.878 K, -0.860 K, and -0.863 K in the second step on the three days. Mean response to the second step is -0.867 K.

The North Cavity temperature changed -0.903 K, -0.933 K, and -0.910 K in the second step on the three days. Mean response to the second step is -0.915 K.


Code and Data


Edit Fri Sep 6 12:01:49 2019 anchal in RED above.

Found out that there is a bored hole in the spacer so the laser doesn't travel in fused silica actually. So the refractive index used above is wrong.

Updated calculations and plots for the same.

Attachment 1: VacCanStepResponseTempAxis.pdf  372 kB  Uploaded Fri Sep 6 13:14:14 2019  | Hide | Hide all
VacCanStepResponseTempAxis.pdf VacCanStepResponseTempAxis.pdf VacCanStepResponseTempAxis.pdf
ELOG V3.1.3-